zarkam21 (zarkam21):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair. x+y=12 -x=-y-10

DivineSolar (divinesolar):

@zarkam21 I want you to participate with me okay?

OpenStudy (girllover123):

@ logan13

DivineSolar (divinesolar):

Girl lover i have this.

OpenStudy (girllover123):

Okay thanks tho

OpenStudy (girllover123):

here is an medal

DivineSolar (divinesolar):

Let me know what you get @zarkam21

DivineSolar (divinesolar):

From the png.

zarkam21 (zarkam21):

Umm I'm stuck

DivineSolar (divinesolar):

How are you stuck :). I provided a clear way for the first step?

zarkam21 (zarkam21):

Is one of the ordered pair -1

DivineSolar (divinesolar):

I will not give out direct answer's as that is a clear violation of the coc.

DivineSolar (divinesolar):

Also -1 is wrong anyways ^_^.

zarkam21 (zarkam21):

Okay Do I solve x+y+-y=12+-y

DivineSolar (divinesolar):

Try this again.

DivineSolar (divinesolar):

Let me know what you get.

zarkam21 (zarkam21):

x=12 ??

DivineSolar (divinesolar):

No, how did you get that?

DivineSolar (divinesolar):

Let me know what you get.

zarkam21 (zarkam21):

Sooo -xy+12=-y-10 ?? :/

DivineSolar (divinesolar):

Now try and solve.

DivineSolar (divinesolar):

After that you simplify both sides of the equation. Then let me know what you get.

zarkam21 (zarkam21):

y+12=-y-10 ??

DivineSolar (divinesolar):

No it would be, y −12= −y − 10

DivineSolar (divinesolar):

Now add y to both sides of that.

zarkam21 (zarkam21):

How did you get -12

zarkam21 (zarkam21):

2y-12=-10

DivineSolar (divinesolar):

it is actually just - 12 not negative 12, and yes, now add 12 to both sides of that equation.

zarkam21 (zarkam21):

y=1 !!

DivineSolar (divinesolar):

Nice :). Now let us do x.

DivineSolar (divinesolar):

Substitute (1) for y in x = −y + 12:

DivineSolar (divinesolar):

Let me know what you get ^-^.

zarkam21 (zarkam21):

-11

DivineSolar (divinesolar):

So where did the negative come from? because when a negative is being added to a positive it just subtracts but does not transfer over the negative.

zarkam21 (zarkam21):

so just 11

DivineSolar (divinesolar):

Correct! So what is your answer/answers?

zarkam21 (zarkam21):

(11,1)

DivineSolar (divinesolar):

Correct :).

zarkam21 (zarkam21):

i mean (1,11)

DivineSolar (divinesolar):

Y is 1, not 11, you had it right the first time :).

DivineSolar (divinesolar):

(11,1) is the correct version.

zarkam21 (zarkam21):

Are you sure

DivineSolar (divinesolar):

Why yes i am ^-^.

zarkam21 (zarkam21):

Thank you so much

DivineSolar (divinesolar):

No problem, anytime!

DivineSolar (divinesolar):

Make sure to close this question ^-^.